Problem 2. Suppose that business travelers and vacationers have the following demand for airline tickets from Karachi to Dubai:

 

Price 

Quantity Demanded

(Business Travelers)

Quantity Demanded

(vacationers)

$150

2100 Tickets

1000 Tickets

$200

2000

 800

$250

$300

1900

1800

 600

 400

 

 

 

 

 

a.   As the price of tickets rises from $150 to $200, what is the price elasticity of demand for (i) business travelers and (ii) vacationers? (Use the midpoint method in your calculations.)

b.    Why might vacationers have a different elasticity from business travelers?

c.    What would you recommend based on the above information for airline to increase the revenue?


1
Expert's answer
2021-05-06T14:03:06-0400


Elasticity of demand depicts how much a consumer responds to change in price level.

(a)

"Ed=\frac{\frac{Q_2-Q_1}{\frac{Q_2+Q_1}{2}}}{\frac{P_2-P_1}{\frac{P_2+P_1}{2}}}"


(i)

For business travelers


"Ped,Ed=\frac{ \frac{2000-2100}{\frac{2000+2100}{2}}}{\frac{200-150}{\frac{200+150}{2}}}"


"Ed=-0.17"

"|Ed|=0.17<1; inelastic"



(ii)

"Ped,Ed=\frac{ \frac{800-1000}{\frac{800+1000}{2}}}{\frac{200-150}{\frac{200+150}{2}}}"



"Ed=-0.78"

"|Ed|=0.78<1;inelastic"


(b)

Vacationers are casual with their plan whereas business travelers are traveling with business purpose thus their elasticity of demand for tickets is less affected with change in price.


(c)

The elasticity of demand curve is inelastic, thus airline can increase price in order to increase Total Revenue